OpenStudy (anonymous):

how to add: (x-1)+(x+1)

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

show work??

OpenStudy (anonymous):

x+x = 2x -1+1 = 0 2x + 0 = 2x

OpenStudy (anonymous):

2x

Parth (parthkohli):

Annoying when people answer directly.

Parth (parthkohli):

Remove the brackets. They do nothing here. \( \color{Black}{\Rightarrow x - 1 + x + 1}\) Add x + x and -1 + 1.

OpenStudy (anonymous):

the brackets aren't needed. might be easier to see as: x-1+x+1

Parth (parthkohli):

\( \color{Black}{\Rightarrow x + x +\cancel{ 1 - 1}}\)

OpenStudy (anonymous):

hey the answer is already given with steps

mathslover (mathslover):

first of all i will say that Please dont answer the questions of the askers directly .. You may get reported or banned http://openstudy.com/code-of-conduct Don't devalue the question/answer process! Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in!

Parth (parthkohli):

Hmm. What is x + x? Well...

OpenStudy (lgbasallote):

want to let @jk_16 talk sometime?

Parth (parthkohli):

Answer must not be given.

OpenStudy (lgbasallote):

there's about 10 replies crammed here =)) let the kid talk lol

mathslover (mathslover):

Hope you all understood the openstudy ...code of conduct .. I am not going to mention his/her name but please help the asker by giving the hints not solution